OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (freckles):

So v=s'

OpenStudy (freckles):

can you find s'?

OpenStudy (freckles):

i'm guessing you need to do this by the definition and no short cuts

OpenStudy (freckles):

?

OpenStudy (freckles):

though you know the derivative=slope

OpenStudy (freckles):

and s is just a line

OpenStudy (anonymous):

is s = -17?

OpenStudy (freckles):

you don't need to find the position at t=1

OpenStudy (freckles):

you need to find the instantaneous velocity at t=1

OpenStudy (anonymous):

so how would I do that exactly?

OpenStudy (freckles):

well i was asking you if you need to use the definition of derivative or could you use short cuts or could you just notice s is a line and and then notice the slope of the line

OpenStudy (jhannybean):

\[v(t) =\frac{d}{dx}[s(t)]\]

OpenStudy (jhannybean):

\[v(1) = s'(1)\]

OpenStudy (anonymous):

im just unsure on how to do this...

OpenStudy (jhannybean):

Take the derivative of \(s(t)\)

OpenStudy (freckles):

err can you tell me how you are allowed to find the derivative? by definition by short cut ways by just noticing it is a line and the slope of y=mx+b is m so therefore the derivative is m

OpenStudy (freckles):

by the way if we have s(t)=mt+b then s'(t)=m for all t

OpenStudy (freckles):

that is what i'm saying with third third way there

OpenStudy (anonymous):

well any way you could show me tbh. my lesson is really confusing and it doesn't make any sense to me

OpenStudy (freckles):

ok well you aren't really telling me what way you are suppose to find the derivative but i actually already did if you realize that s(t)=-8-9t is in the form s(t)=mt+b

OpenStudy (jhannybean):

I would start by understanding what the slope of a line represents in calculus terms. Did you learn this yet?

OpenStudy (anonymous):

well then is it just -9?

OpenStudy (freckles):

yep the slope of y=-9t-8 is -9 therefore the derivative of y is y'=-9

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @freckles by the way if we have s(t)=mt+b then s'(t)=m for all t \(\color{blue}{\text{End of Quote}}\) That's what was said :)

OpenStudy (freckles):

But I'm kinda wondering since this was an easy function if the directions were to find the derivative by using the formal definition of derivative

OpenStudy (jhannybean):

You could do that too.

OpenStudy (jhannybean):

\[\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

well could you try and show me the other way as well?:)

OpenStudy (jhannybean):

Ok, \[ s(t) = -8 - 9t\] Find : 1. \(s(x+h)\) 2. \(s(x)\)

OpenStudy (jhannybean):

Instead of the t use h or convenience, it just works to fit the definition, HAH

OpenStudy (anonymous):

what would be h? just 1?

OpenStudy (jhannybean):

\[s(x+h) = -8-9(x+h)\]\[s(x) = -8-9(x)\]

OpenStudy (jhannybean):

h is a representation for a small change in the value of your function.

OpenStudy (jhannybean):

so \[\lim_{x\rightarrow 0} \frac{-8-9x-9h-(-8-9x)}{h}\]\[\lim_{x\rightarrow 0}\frac{-9h}{h} = -9\]

OpenStudy (anonymous):

ok, where do I go from there? you see they give us some equation f(x+h)-f(x)/h = lim h->0, which doesn't make sense

OpenStudy (anonymous):

lol wait you just showed me:p sorry im having connection problems with this website rn and idk why

OpenStudy (jhannybean):

I've got to go, hopefully @freckles can provide you with a better explanation than me :P Good luck!

OpenStudy (anonymous):

ok so h will cancel out and with the x plugged in as 1, the answer will equal -9?:)

OpenStudy (freckles):

\[v(t)=s'(t)=\lim_{h \rightarrow 0}\frac{s(t+h)-s(t)}{h}\] h is a small change as said above we want this small change to approach 0 so our slope of our secant lines from (t+h,s(t+h)) to (t,s(t)) will approach the slope of the tangent line

OpenStudy (freckles):

\[s(t)=-8-9t \\ \text{ so we can find } s(t+h) \text{ by replacing the } t \text{ above with } t+h \\ s(t+h)=-8-9(t+h)\] inputting into our formula: \[v(t)=\lim_{h \rightarrow 0} \frac{[-8-9(t+h)]-[-8-9t]}{h}\] simplify the top as @Jhannybean did above

OpenStudy (freckles):

then you will be allowed to rid yourself of the h on bottom by cancelling a common factor of h on top with the one on bottom and therefore you will be able to plug in 0 for h after that if any h's remain that is

OpenStudy (anonymous):

ahh ok thanks so much:)))

OpenStudy (freckles):

there is also another form of definition you can use \[s'(t)=\lim_{z \rightarrow t} \frac{s(z)-s(t)}{z-t}\] that from right there is probably what you have more commonly associated slope with \[s'(t)=\lim_{z \rightarrow t}\frac{[-8-9z]-[-8-9t]}{z-t} \\ =\lim_{z \rightarrow t}\frac{-9z+9t}{z-t} =\lim_{z \rightarrow t}-9 \frac{z-t}{z-t}=-9(1)=-9\]

OpenStudy (freckles):

these forms are equivalent you can show they are equivalent by substitution

OpenStudy (freckles):

let h=z-t if z->t then h->z-z=0 so h->0 \[s'(t)=\lim_{z \rightarrow t}\frac{s(z)-s(t)}{z-t} \\= \lim_{h \rightarrow 0}\frac{s(t+h)-s(t)}{h}\] since if h=z-t then h+t=z